the sum of 2 numbers is -23, and the difference of the larger and smaller is 81. What are the two numbers?
Answer by checkley77(12844)

X+Y=-23
X-Y=81 ADD.
----------------
2X=58
X=58/2
X=29 ANS FOR THE LARGER NUMBER.
29+Y=-23
Y=-23-29
Y=-52 ANS.
PROOF:
29-(-52)=81
29+52=81
81=81